OpenStudy (anonymous):

Starting with the graph of f(x)=4^x , write the equation of the graph that results from (a) shifting f(x) 4 units upward. y=? (b) shifting f(x) 9 units to the right. y? c) shifting f(x) about the y-axis. y=?

OpenStudy (anonymous):

a) y = 4^x + 4 b) y= 4^(x-9)

OpenStudy (anonymous):

not sure about c - ill check that out

OpenStudy (anonymous):

Thanks I got c it's 1/4^x

OpenStudy (anonymous):

yes - correct

OpenStudy (anonymous):

yw